# Improved implicit diffusion model with knowledge distillation to estimate the spatial distribution density of carbon stock in remote sensing imagery

### v1

@@ Abstract
The forest serves as the most significant terrestrial carbon stock mechanism, effectively reducing atmospheric CO2 concentrations and mitigating climate change. Remote sensing provides high data accuracy and enables large-scale observations. Optical images facilitate long-term monitoring, which is crucial for future carbon stock estimation studies. This study focuses on Huize County, Qujing City, Yunnan Province, China, utilizing GF-1 WFV satellite imagery. The KD-VGG and KD-UNet modules were introduced for initial feature extraction, and the improved implicit diffusion model (IIDM) was proposed. The results showed: (1) The VGG module improved initial feature extraction, improving accuracy, and reducing inference time with optimized model parameters. (2) The Cross-attention + MLPs module enabled effective feature fusion, establishing critical relationships between global and local features, achieving high-accuracy estimation. (3) The IIDM model, a novel contribution, demonstrated the highest estimation accuracy with an RMSE of 12.17%, significantly improving by 41.69% to 42.33% compared to the regression model. In carbon stock estimation, the generative model excelled in extracting deeper features, significantly outperforming other models, demonstrating the feasibility of AI-generated content in quantitative remote sensing. The 16-meter resolution estimates provide a robust basis for tailoring forest carbon sink regulations, enhancing regional carbon stock management.
@@
Introduction
@@
Given the increasing threat of global climate change, human activities pose unprecedented challenges to Earth’s ecosystems [1]. As a result, forest resource management and use have become primary concerns for researchers and policymakers worldwide [2]. Forest ecosystems provide a variety of ecological services and play a crucial role as regulators of the global carbon cycle, contributing significantly to climate change mitigation [3]. It is imperative to monitor and quantify the temporal and spatial variations in regional carbon stocks to understand the dynamics of forest carbon sources and sinks and to develop effective regulatory strategies [4, 5].
Forest carbon stock monitoring is increasingly integrating ground-based sample data with satellite remote sensing observations [6, 7, 8, 9]. While ground monitoring provides precise data, it is hindered by time-consuming and labor-intensive processes, making it impractical for large-scale observations [10, 11]. In contrast, remote sensing inversion methods offer a viable solution by overcoming the limitations of sample-based monitoring, enabling more efficient and accurate monitoring with continuous improvements [12, 13]. Spectral information-based methods can infer vegetation growth status and identify vegetation types within ecosystems, aiding in estimating carbon stocks. However, despite their high accuracy, these methods often struggle to explore intricate nonlinear relationships due to image quality issues, resulting in limited improvements in estimation accuracy [14].
Methods based on structural information can directly measure biomass and carbon stock, but their effectiveness is constrained by the resolution and coverage of remote sensing images [10]. Physical model-based approaches estimate the carbon stock by developing carbon-cycle models that simulate the carbon sink and cycle processes within ecosystems. Although this method can account for the variations and complexities of different ecosystems, it requires precise ecological parameters and robust data support [15]. Conversely, machine learning-based methods uncover deeper relationships between image data and carbon stock, enabling rapid and efficient estimates [16, 17]. Integrating sample monitoring data with remote sensing data to establish a universal and high-precision remote sensing monitoring model for regional forest carbon stock has become an urgent priority.
In summary, we have successfully reframed the carbon stock estimation problem as an image translation task and utilized the implicit diffusion model as our foundational network structure. Using GF-1 WFV satellite imagery as the primary data source and adapting our methodology to the specific characteristics of Huize County, we achieved precise estimations of the regional carbon stock distribution density. This work provides a critical theoretical foundation for informing forest carbon sink regulation strategies and decision-making. Our main contributions include:
  • The VGG module, enhanced through knowledge distillation, was incorporated into the implicit diffusion model for initial feature extraction and to construct the encoder-decoder structure. Similarly, the UNet model was distilled, resulting in a reduction in the number of parameters.
  • Attention mechanisms combined with MLP were added for feature fusion, enabling the capture of relationships between local and global features and achieving high-accuracy estimates of carbon stock density distribution.
  • The improved implicit diffusion model (IIDM) reduced inference time and enhanced accuracy, with experimental results demonstrating the feasibility of artificial intelligence-generated content (AIGC) in quantitative remote sensing.
Related work
Key issues in carbon stock estimation
@@
Data selection. Ground monitoring, while providing precise data, is impractical for large-scale observations due to its labor-intensive and time-consuming nature [10, 11]. Remote sensing inversion methods mitigate these issues, improving both efficiency and accuracy [12, 13]. Among the most promising remote sensing resources for additional data sets are medium-resolution (10$\sim$30 m) optical data. The long operational lifetimes of these satellites make them ideal for continuous monitoring of forest dynamics [18]. When utilizing multispectral images as data sources, carbon stock calculations have mainly focused on estimating biomass and stock, with relatively few studies directly estimating carbon stock from image data.
Method selection. Zhang et al. [19] combined ground observations, MODIS, GLAS, climate, and terrain data with the random forest algorithm to generate a 1 km resolution aboveground biomass map for China, achieving an interpretability of 75% and an RMSE of 45.5 Mg/ha. Puliti et al. [18] estimated the total change in forest aboveground biomass over approximately 1.4 million hectares in northern Norway using National Forest Inventory (NFI), Sentinel-2, and Landsat data. Chopping et al. [20] estimated aboveground biomass in the southwestern U.S. from 2000 to 2015 using the Multi-angle Imaging Spectro Radiometer (MISR), achieving an RMSE of 37.0 Mg/ha. Despite the detailed spatial features and rich surface texture information provided by medium-resolution remote sensing images, these images typically have limited spectral bands, and existing carbon stock data products usually exhibit low resolution. Therefore, the key to accurately estimating carbon sinks lies in developing effective algorithms capable of extracting deep-seated features.
Challenges for quantitative remote sensing
One of the most significant challenges in quantitative remote sensing is effectively managing complex non-linear relationships to improve data accuracy [21, 22]. Although remote sensing inversion methods can efficiently monitor and estimate forest carbon stocks, they still face obstacles such as poor image quality and complex identification of vegetation growth and types. Consequently, there is a need for the development of new techniques and methods in quantitative remote sensing to address these issues and improve the accuracy and efficiency of data processing and estimation [23, 21].
Deep learning surpasses traditional machine learning methods in data processing capabilities, automated feature learning, and generalization, making it well-suited for handling complex, high-dimensional big data and achieving superior performance in various tasks [24, 21]. This paper investigates the feasibility of integrating quantitative remote sensing with computer vision, using generative modeling in deep learning as a case study. We explore how combining these fields can enhance data processing and analysis in quantitative remote sensing.
The Intersection of Quantitative Remote Sensing and Computer Vision
The widespread adoption of deep learning techniques has enabled the extraction of deep-seated features to their fullest potential, leading to significant improvements in accuracy [10, 25, 26].
Machine learning. Many current studies have utilized traditional machine learning methods such as Ordinary Least Squares (OLS), Random Forest (RF), and Support Vector Regression (SVR) [27, 21, 24]. While these methods are straightforward and efficient, their estimation accuracy is often limited. The application of deep learning methods in carbon stock estimation remains relatively unexplored. Deep learning models, particularly generative models, have the ability to generate new samples, enhance data utilization, improve model generalization, and address issues related to insufficient or incomplete data, making them well-suited for carbon stock estimation. Generative models aim to model the distribution of real data by learning from multi-modal data using implicit variable techniques.
Generative model. The Generative Adversarial Network (GAN), proposed by Goodfellow et al. [28], comprises generators and discriminators that engage in adversarial training to produce high-quality samples. However, GAN training can be unstable and prone to mode collapse, requiring careful parameter tuning and attention throughout the training process. Variational Autoencoders (VAE), proposed by Kingma and Welling [29], enhance sample continuity and controllability by learning a continuous latent space. Despite this, VAEs often underperform compared to GANs in generating high-quality, realistic samples, possibly due to the vagueness inherent in their latent space.
Diffusion model. The diffusion model, first introduced in 2015 [30], aims to progressively remove Gaussian noise from training images, functioning similarly to a series of denoising autoencoders. Variants like the Latent Diffusion Model (LDM) train autoencoders to map images into low-dimensional latent spaces. The diffusion model excels in generating high-quality samples with flexibility, stability during training, and resistance to mode collapse, effectively bridging the gaps between GANs and VAEs by combining their strengths. Subsequent models such as DDPM [31], DDIM [32], and Stable Diffusion [33] have extended the diffusion model’s applications to image generation, consistently achieving state-of-the-art (SOTA) results. These models have been widely applied in image translation tasks. Building on these advancements, we applied the diffusion model structure to estimate carbon stock in remote sensing images, aiming to achieve improved accuracy in this domain.
Materials
Study area
@@
Huize County is part of Qujing City, Yunnan Province, China, located at the intersection of Yunnan, Sichuan, and Guizhou Provinces. Covering an area of 5,889 km2 (Figure 1), Huize features a ladder-shaped terrain with higher peaks in the west and lower valleys in the east. The county experiences a typical temperate plateau monsoon climate, with climates ranging from subtropical in the south to cold temperate in the north. Elevations vary widely, from the highest peak at 4,017 m to the lowest point at 695 m. According to Forest Management Inventory data, Huize County has approximately 3,080.53 km2 of forest land, with arbor forest land covering approximately 2,538.20 km2 (82.39%). Rich in forest resources, Huize provides diverse research samples, although its complex landscape poses significant challenges.
@@
Data sources
@@
Survey data. The study utilized 2020 data from the Forest Management Inventory obtained from the Huize County Forestry and Grass Bureau. This comprehensive data includes over 70 attributes such as forest area and type, providing an objective overview of forest resources. It aids in understanding and evaluating the status, trends, distribution, and ecological environment of forest resources, providing a scientific basis for forestry management, protection, and sustainable utilization.
@@
Satellite imagery. The study primarily used GF-1 WFV image data. Launched by China in 2013, the GF-1 satellite provides medium-resolution remote sensing with a spatial resolution of 16 meters. It is widely used in agriculture, forestry, environmental monitoring, and more. For time consistency, images from August 27, 2020, were used for model training, covering the entire study area. Data details are provided in Table A1.
@@
Elevation data. The study utilized ALOS PALSAR DEM data with a spatial resolution of 12.5 meters, achieving full coverage of the study area with four-view images (Table A1). ALOS PALSAR, a satellite equipped with a Synthetic Aperture Radar (SAR) sensor, scanned the Earth’s surface with radar beams to measure topographic altitude and surface characteristics. The L-band SAR of PALSAR provided detailed, all-weather interferometric measurements from 2006 to 2011.
@@
Canopy height. The study utilized a 10-meter vegetation canopy height dataset from the ETH Global Sentinel-2 [34]. Lang et al. [34] fused GEDI and Sentinel-2 data to develop a probabilistic deep learning model for retrieving canopy height from Sentinel-2 imagery globally. The global canopy height map is based on Sentinel-2 images taken between May and September 2020. We extracted the relevant data for Huize County.
Methods
Pre-processing
Pre-processing procedures are described in Appendix Section 1. The pre-processing process includes carbon stock & distribution density calculation, and forest/non-forest areas identification.
Knowledge distillation (KD) module

### v2

# IIDM: Improved Implicit Diffusion Model with Knowledge Distillation to Estimate the Spatial Distribution Density of Carbon Stock in Remote Sensing Imagery

@@ Abstract
Forests play a critical role as terrestrial carbon sinks, effectively reducing atmospheric CO2 concentrations and mitigating climate change. Remote sensing facilitates large-scale, high-accuracy carbon stock estimation, with optical imagery enabling long-term monitoring. This study focuses on Huize County, Qujing City, Yunnan Province, China, utilizing GF-1 WFV satellite imagery. We propose an Improved Implicit Diffusion Model (IIDM), integrating Knowledge Distillation (KD) for feature extraction and spatial distribution modeling. The key contributions are as follows: (1) A knowledge distillation-enhanced VGG (KD-VGG) module for efficient feature extraction, reducing inference time while preserving crucial spectral and spatial information. (2) A KD-UNet architecture with cross-attention and MLP modules for improved feature fusion, effectively capturing global-local relationships and enhancing estimation accuracy. (3) A novel IIDM framework incorporating generative diffusion modeling and implicit neural representations, enabling high-resolution spatial carbon stock estimation. IIDM achieves state-of-the-art accuracy, with an RMSE of 12.17%, outperforming traditional regression models by 41.69% to 42.33%. The model effectively extracts deep spatial and spectral features, surpassing existing methods and reinforcing the feasibility of AI-generated content in quantitative remote sensing. The high-resolution (16-meter) carbon stock distribution maps provide a solid foundation for refining forest carbon sink regulations and improving regional carbon stock management strategies.
@@
Carbon stock , GF-1 WFV , Distribution density , Diffusion model , Knowledge distillation
@@
[label1]organization=Faculty of Computer Science and Information Technology,addressline=Universiti Malaya, city=Kuala Lumpur, postcode=50603, state=Kuala Lumpur, country=Malaysia
[…]
[label2]organization=Faculty of Geography and Remote Sensing,addressline=Guangzhou University, city=Guangzhou, postcode=510006, state=Guangdong Province, country=China
{highlights}
KD-VGG and KD-UNet models significantly improve feature extraction accuracy and reduce inference time through knowledge distillation.
Cross-attention + MLPs module effectively fuses global and local features, leading to superior carbon stock estimation accuracy.
Improved Implicit Diffusion Model (IIDM) achieves 41.69%-42.33% higher accuracy than traditional regression models, with an RMSE of 12.17%.
1 Introduction
@@
Given the increasing threat of global climate change, human activities pose unprecedented challenges to Earth’s ecosystems Lenoir et al. (2020). As a result, forest resource management and utilization have become primary concerns for researchers and policymakers worldwide Pecchi et al. (2019). Forest ecosystems play a pivotal role in regulating the global carbon cycle, offering critical ecological services and contributing significantly to climate change mitigation Brando et al. (2019). Accurately estimating the spatial distribution and temporal variation of regional carbon stocks is essential for understanding forest carbon dynamics and formulating effective regulatory strategies Nolan et al. (2021); Hua et al. (2022).
Recent advancements in remote sensing technology have facilitated large-scale forest carbon stock monitoring by integrating ground-based sample data with satellite observations Hurtt et al. (2019); Sun and Liu (2020); Lee et al. (2021); Santoro et al. (2022). While ground surveys provide high-precision data, they are labor-intensive, time-consuming, and impractical for large-scale monitoring Xiao et al. (2019); Gao et al. (2023a). In contrast, remote sensing-based inversion models provide efficient and scalable solutions for estimating carbon stocks across large regions Long et al. (2020); Gray et al. (2020). Conventional spectral information-based methods infer vegetation conditions from optical imagery to estimate carbon stocks, but they struggle to model complex nonlinear relationships due to variations in image quality, spectral distortions, and spatial resolution limitations Chen et al. (2019).
Structural information-based methods measure forest biomass and carbon stock more directly, but their effectiveness is constrained by the spatial resolution and coverage of remote sensing images Xiao et al. (2019). Physical model-based approaches utilize ecosystem carbon-cycle simulations, offering theoretical insights but requiring extensive field-measured ecological parameters, making them challenging to generalize for large-scale applications He et al. (2021). Machine learning (ML)-based models have emerged as a powerful alternative, revealing hidden patterns in remote sensing data and improving estimation accuracy without reliance on predefined physical assumptions Odebiri et al. (2021); Wang et al. (2022). However, traditional ML models such as regression and decision tree-based algorithms often fail to fully capture the spatial dependencies and hierarchical feature relationships within remote sensing images.
To address these limitations, we introduce the Improved Implicit Diffusion Model (IIDM), a novel framework that integrates diffusion models with implicit neural representations for enhanced carbon stock estimation. IIDM leverages a generative approach to refine spatial distribution density estimation, surpassing conventional ML techniques in extracting deep spatial features. By incorporating knowledge distillation, our method optimizes feature extraction while reducing model complexity, thereby improving both efficiency and accuracy.
This study applies IIDM to Huize County, using GF-1 WFV satellite imagery as the primary data source to achieve high-precision regional carbon stock estimations. Our approach provides a theoretical foundation for improved forest carbon sink regulations and data-driven decision-making in climate change mitigation. The main contributions of our work are as follows:
  • 1.
    We integrate knowledge distillation into the VGG module within the diffusion model, enhancing initial feature extraction. Additionally, UNet undergoes distillation, leading to a significant reduction in parameters while preserving feature representation, thereby improving inference efficiency.
  • 2.
    We introduce a cross-attention mechanism combined with MLPs to enhance feature fusion, effectively capturing complex global-local dependencies and improving the spatial distribution accuracy of carbon stock density estimation.
  • 3.
    We propose the IIDM framework, which optimizes the inference process while achieving state-of-the-art estimation accuracy. This demonstrates the feasibility of leveraging artificial intelligence-generated content (AIGC) for high-precision remote sensing applications.
2 Related Work
2.1 Key Issues in Carbon Stock Estimation
Definition and Importance. Carbon stock refers to the total amount of carbon stored in a forest ecosystem, including vegetation, soil, and biomass. It plays a crucial role in global climate regulation by acting as a carbon sink that absorbs atmospheric CO2. However, estimating carbon stock with high spatial resolution remains challenging due to environmental heterogeneity and the limitations of existing remote sensing techniques. Understanding the spatial distribution density of carbon stock is essential for precise forest resource management, carbon trading, and climate change mitigation policies.
@@ 2.1 Key Issues in Carbon Stock Estimation
Data Selection. Ground monitoring provides precise data but is impractical for large-scale observations due to its labor-intensive and time-consuming nature Xiao et al. (2019); Gao et al. (2023a). Remote sensing inversion methods mitigate these issues, improving both efficiency and accuracy Long et al. (2020); Gray et al. (2020). Among the most promising remote sensing resources for additional datasets are medium-resolution (10$\sim$30 m) optical data. The long operational lifetimes of these satellites make them ideal for continuous monitoring of forest dynamics Puliti et al. (2021). However, optical multispectral data, though widely used, suffer from spectral band limitations that hinder their ability to capture fine-scale variations in carbon stock. This challenge necessitates the adoption of deep learning models capable of extracting hierarchical feature representations to improve estimation accuracy.
Method Selection. Traditional carbon stock estimation methods integrate field observations with remote sensing imagery. Zhang et al. Zhang et al. (2019) combined MODIS, GLAS, climate, and terrain data with random forests to produce a 1 km resolution biomass map for China, achieving an interpretability of 75% and an RMSE of 45.5 Mg/ha. Puliti et al. Puliti et al. (2021) estimated forest biomass change over 1.4 million hectares in Norway using Sentinel-2 and Landsat data. Chopping et al. Chopping et al. (2022) used MISR to estimate biomass in the southwestern U.S. from 2000 to 2015, achieving an RMSE of 37.0 Mg/ha. While these methods offer valuable insights, their reliance on traditional machine learning models limits their ability to capture intricate spatial dependencies and nonlinear feature relationships critical for precise carbon stock estimation.
To overcome these challenges, we introduce the Improved Implicit Diffusion Model (IIDM), a generative deep learning approach that enhances carbon stock estimation by leveraging feature learning techniques, knowledge distillation, and multi-scale spatial feature extraction.
2.2 Challenges in Quantitative Remote Sensing
One of the most significant challenges in quantitative remote sensing is effectively managing complex non-linear relationships while improving estimation accuracy Han et al. (2023); Yasir et al. (2023). Although remote sensing inversion methods allow efficient monitoring of forest carbon stocks, obstacles such as low image quality, spectral noise, and environmental heterogeneity still hinder their performance. Current methods struggle to generalize across diverse ecological conditions, necessitating advanced deep learning models capable of dynamically adapting to different environmental settings.
This study introduces IIDM, which integrates implicit neural representations into the diffusion framework to refine carbon stock estimation, effectively addressing issues related to image degradation, data sparsity, and inconsistent spectral responses.
Deep learning surpasses traditional machine learning methods in handling high-dimensional data and extracting hierarchical representations, making it particularly effective for remote sensing applications Wang et al. (2023); Han et al. (2023). Unlike traditional models that rely on manually selected features, IIDM employs an end-to-end learning paradigm that automatically refines spatial feature representations, improving accuracy and robustness in carbon stock estimation.
2.3 The Intersection of Quantitative Remote Sensing and Computer Vision
Recent advancements in deep learning have significantly enhanced feature extraction in remote sensing applications Xiao et al. (2019); Matinfar et al. (2021); Pham et al. (2023). In particular, the fusion of deep learning and artificial intelligence-generated content (AIGC) has introduced a new paradigm for high-precision carbon stock estimation. This approach reconstructs spatial distribution density with enhanced accuracy, leveraging generative models to synthesize missing information and refine predictions. Our proposed IIDM follows this paradigm by employing a diffusion-based generative framework to enhance spatial feature representation and estimation accuracy.
Machine Learning. Many current studies have utilized traditional machine learning methods such as Ordinary Least Squares (OLS), Random Forest (RF), and Support Vector Regression (SVR) for carbon stock estimation Li et al. (2024); Han et al. (2023); Wang et al. (2023). While effective for linear and moderately nonlinear relationships, these models often fail to generalize in heterogeneous landscapes. Unlike these traditional approaches, IIDM leverages deep feature learning to model complex nonlinear dependencies between remote sensing data and carbon stock distribution. Furthermore, IIDM benefits from the knowledge distillation framework, which significantly improves efficiency while maintaining feature representational power, addressing the high computational demands often associated with deep learning-based estimation models.
Generative Models. The Generative Adversarial Network (GAN), proposed by Goodfellow et al. Goodfellow et al. (2014), employs adversarial training to synthesize realistic samples. However, GAN training is inherently unstable, prone to mode collapse, and requires extensive hyperparameter tuning. Variational Autoencoders (VAE), introduced by Kingma and Welling Kingma and Welling (2013), improve sample diversity but often produce blurry reconstructions. IIDM overcomes these limitations by utilizing an implicit diffusion framework, which refines feature representations progressively, ensuring stable and accurate carbon stock estimations. Unlike GANs, which require adversarial optimization, and VAEs, which struggle with sample sharpness, diffusion models provide a principled approach to generating high-quality feature distributions.
Diffusion Models. The diffusion model, first introduced by Sohl-Dickstein et al. Sohl-Dickstein et al. (2015), progressively removes Gaussian noise from training images, functioning similarly to denoising autoencoders. Latent Diffusion Models (LDMs) refine this approach by mapping images into compact latent spaces. Unlike GANs, diffusion models offer stable training, improved sample diversity, and resilience to mode collapse. Our proposed IIDM builds upon the strengths of diffusion models while addressing their computational inefficiencies through knowledge distillation and multi-scale feature extraction. Traditional diffusion models, such as DDPM Ho et al. (2020), DDIM Song et al. (2020), and Stable Diffusion Rombach et al. (2022), focus on generating high-quality image samples, whereas IIDM is explicitly designed for structured regression tasks, such as carbon stock estimation, by integrating deep hierarchical feature representations and an optimized implicit neural framework.
In summary, IIDM bridges the gap between conventional machine learning, generative models, and diffusion models by providing a novel knowledge distillation-enhanced diffusion framework tailored for high-precision spatial carbon stock estimation. This represents a significant step forward in AI-driven remote sensing and quantitative environmental modeling.
3 Materials
3.1 Study area
@@
Huize County, located in Qujing City, Yunnan Province, China, covers an area of 5,889 km2 (Figure 1). The county features a ladder-shaped terrain, with elevations ranging from the highest peak at 4,017 m to the lowest point at 695 m. This complex topography significantly influences forest carbon distribution and presents challenges for remote sensing-based carbon stock estimation. The region experiences a temperate plateau monsoon climate, transitioning from a subtropical climate in the south to a cold temperate climate in the north, which further impacts vegetation growth patterns and carbon sequestration capacity. According to Forest Management Inventory data, Huize County has approximately 3,080.53 km2 of forest land, with arbor forest land covering approximately 2,538.20 km2 (82.39%). The rich forest resources and diverse ecological conditions make Huize County an ideal site for carbon stock estimation research, despite the challenges posed by its complex geographical landscape.
@@
3.2 Data sources
This study integrates multiple data sources, including field survey data, satellite imagery, and topographic datasets, to enhance carbon stock estimation accuracy.
@@ 3.2 Data sources
Survey data. The study utilized 2020 data from the Forest Management Inventory obtained from the Huize County Forestry and Grass Bureau. This dataset comprises over 70 attributes, such as forest area and type, enabling a comprehensive analysis of forest distribution and carbon storage dynamics. These data provide a scientific basis for forestry management, protection, and sustainable utilization.
@@
Satellite imagery. The study primarily used GF-1 WFV image data. Launched by China in 2013, the GF-1 satellite provides medium-resolution remote sensing with a spatial resolution of 16 meters. It features four spectral bands (red, green, blue, and near-infrared), which are essential for vegetation monitoring and carbon stock estimation. To ensure time consistency, images from August 27, 2020, were used for model training, covering the entire study area. Data details are provided in Table A1.
@@
Elevation data. The study utilized ALOS PALSAR DEM data with a spatial resolution of 12.5 meters, achieving full coverage of the study area with four-view images (Table A1). ALOS PALSAR, a satellite equipped with a Synthetic Aperture Radar (SAR) sensor, provides detailed, all-weather topographic measurements, which are crucial for understanding terrain-induced variations in carbon stock distribution.
@@
Canopy height. The study utilized a 10-meter vegetation canopy height dataset from the ETH Global Sentinel-2 Lang et al. (2022). Lang et al. Lang et al. (2022) fused GEDI and Sentinel-2 data to develop a probabilistic deep learning model for retrieving canopy height from Sentinel-2 imagery globally. The global canopy height map is based on Sentinel-2 images taken between May and September 2020. This dataset offers critical insights into forest structure and biomass distribution, supporting high-accuracy carbon stock modeling.
By integrating these diverse datasets, we ensure a balance between spatial coverage and estimation precision, making them well-suited for carbon stock modeling in complex terrains such as Huize County.
4 Methods
4.1 Overview
To estimate carbon stock spatial distribution density with high precision, we propose an Improved Implicit Diffusion Model (IIDM) that integrates knowledge distillation, implicit neural representations, and diffusion models. The workflow consists of several key components (Figure 2):
  1. 1.
    Pre-processing Pipeline. A pre-processing pipeline to extract relevant forest data
  2. 2.
    Knowledge Distillation (KD). A knowledge distillation (KD) module for efficient feature extraction.
  3. 3.
    Denoising Model. A denoising model to refine spatial features.
  4. 4.
    Implicit Representation. An implicit representation module to enhance feature continuity.
  5. 5.
    Diffusion-Based Framework. A diffusion-based generative framework for accurate prediction.
Figure 2: Overview.
4.2 Pre-processing
Pre-processing procedures are described in Appendix Section A. The process includes three key steps: (1) calculating carbon stock based on survey data using IPCC default coefficients, (2) computing carbon stock distribution density by normalizing canopy height weights, and (3) identifying forest and non-forest areas using the F-Pix2Pix model. These steps ensure accurate estimation of carbon stock while filtering irrelevant areas for analysis.
4.3 Knowledge distillation (KD) module